# Interpretable Evaluation of AI-Generated Content with Language-Grounded Sparse Encoders

### v1

@@ Results
Figure 10: CLIP’s blind spots revealed through LanSE’s fine-grained analysis. Examples where CLIP assigns high scores (>0.28>0.28) to images with clear semantic mismatches. LanSE successfully detects unmentioned elements that CLIP overlooks.
@@
Our results reveal a striking pattern in recent generative models: while they achieve strong semantic alignment, often exceeding that of real-world datasets, they consistently struggle with visual realism, physical plausibility, and content diversity. In particular, recent models frequently exhibit distinctive stylistic signatures that deviate from photorealistic representation, generate distorted surfaces, and produce physically implausible structures. Their content diversity remains significantly below that of natural image distributions. These findings suggest the machine learning community should recalibrate its priorities, shifting focus from semantic alignment, a largely solved problem, toward these more nuanced aspects of visual generation.
From an interpretability perspective, our systematic curation of 5309 interpretable LanSE neurons collected from 3.75 million candidates demonstrates that mechanistic interpretability features can reliably signal visual patterns with over 90% accuracy, forming a robust foundation for evaluating generative models. These features enable targeted assessment of specific quality dimensions while providing decomposed evaluation axes between recent models compared with existing metrics. Rather than treating interpretability features as ancillary tools for post-hoc analysis [60], we propose a paradigm shift: positioning them as structural infrastructure for diagnostic assessment. LanSE’s diverse visual patterns, each grounded in natural language explanations, enable evaluation methods that are not merely quantitative but genuinely explanatory, revealing not just failure rates, but specific failure types and their underlying causes.
Beyond evaluation, LanSE opens several avenues for future development. Looking ahead, researchers should explore expanding LanSE’s scope along multiple dimensions. Task-specific neuron groups could address currently unmeasured qualities such as aesthetic appeal, compositional balance, or social bias. The framework could extend to other generative modalities, with video generation evaluation presenting immediate opportunities for temporal consistency and motion physics assessment. We envision integrating LanSE directly into model training pipelines, where generative models could receive fine-grained feedback through interpretable, feature-level signals. Additionally, LanSE’s architecture naturally supports various downstream applications including multi-label explanation generation [61] (See Appendix B), feature-based information retrieval [62], and systematic slice discovery [63]. By establishing language-grounded sparse encoders as foundational infrastructure, we aim to catalyze a broader transformation in machine learning, one that embraces richer, more interpretable feedback throughout the evaluation, training, and deployment lifecycle.
Limitations
While LanSE provides interpretable and actionable evaluation signals, it is subject to several limitations. First, its ability to assess generative outputs is constrained to visual patterns that are describable in natural language, as neuron validation requires human or LMM-readable explanations. This introduces an anthropocentric bias, subtle, cognitively opaque, or sub-symbolic failure modes may evade detection. Second, LanSE’s reliance on large multimodal models (LMMs) for neuron interpretation may introduce inconsistency or bias into the feature categorization process. Finally, the scope and quality of evaluation are bounded by the diversity and granularity of the neuron groups discovered. Despite these constraints, we argue that LanSE remains a practical and human-aligned approach: the failure patterns that matter most to end users tend to be those that are both perceptible and articulable. We summarize this guiding philosophy as: ”LanSE evaluates with what human language can describe. What we care about, we evolve language to describe.” A more detailed discussion of limitations is presented in Appendix I.
@@
In this section, we present the the details of our proposed Language-Grounded Sparse Encoders method (See Figure 11). Our method addresses the challenge of transforming mechanistic interpretability features into practical evaluation infrastructure through a systematic three-stage pipeline. It basically proceeds in three stages: (1) discovering sparse latent features using sparsity-driven interpretability modules (2) categorizing these features automatically with the help of large multimodal models (LMMs), and (3) building single-modality LanSEs through a modality distillation procedure to support the computation of prompt match metrics. We extend our method with a human-involved interactive framework that can support the discovery of targeted types of error patterns in Appendix D, which is used to find the distortion group and the structure group of neurons in the LanSEs we have constructed.
Figure 11: The general pipeline to construct LanSEs with SAE as the feature discovery tool. §: The Visual LanSE and the Textual LanSE are distilled from the Visual-Textual LanSE, ensuring that the three LanSEs have an aligned latent space.
Dataset Augmentation
For datasets, we have used captions from MSCOCO [64], Flickr8k and Flickr30k [65], TextCaps [66], and NoCaps [67]. We augment these image-caption datasets with generated data using a variety of recently developed generative models [54, 55, 56, 57, 68] to get a more diverse range of images with 253,213 images used in total. All of these image-caption datasets are used downstream to construct LanSE models and identify the images that can activate a LanSE neuron.
Automatic Feature Discovery with Sparsity-Driven Interpretability Modules
To discover interpretable patterns, we encode image-caption pairs into embeddings, and use the concatenations of the embeddings to train sparse autoencoders. These sparse autoencoders aim to extract interpretable features from representations with superposition [45], to produce monosemantic neurons [43], latent space neurons that trigger according to a specific concept.
Given the image-caption pairs, $(\mathbf{x_{i}},\mathbf{y_{i}})\sim\mathbb{P}(\mathcal{X},\mathcal{Y}),(i=1,2...N)$, we encode them into a joint representation $\mathbf{e_{i}}\in\mathbb{R}^{d}$ by concatenating the outputs of two fixed pretrained encoders $\mathrm{E_{x}}$ and $\mathrm{E_{y}}$:
$\mathbf{e_{i}}=[\mathrm{E_{x}}(\mathbf{x_{i}}),\ \mathrm{E_{y}}(\mathbf{y_{i}})].$ (1)
Sparse autoencoders are neural networks designed to reconstruct input data while enforcing sparsity in their latent representations, encouraging each neuron to respond to distinct, interpretable features [69]. We subsequently train multiple k-sparse autoencoder [70] each comprised of one encoder $\mathbf{z_{i}}=\mathrm{Enc}(\mathbf{e_{i}})=\mathrm{ReLU}(\mathbf{w}\cdot\mathbf{e_{i}}+\mathbf{b})$ and one decoder $\mathrm{Dec}(\mathbf{z_{i}})=\mathrm{ReLU}(\mathbf{w}\cdot\mathbf{z_{i}}+\mathbf{b})$ to reconstruct $\mathbf{e_{i}}$ from a sparse latent vector, $\mathbf{z_{i}}=\mathrm{Enc}(\mathbf{e_{i}})$. To enforce sparsity, we adopt a top-$k$ activation scheme: for each $\mathbf{z_{i}}$, we retain only the $k$ highest-magnitude activations and set the rest to zero, yielding the following training objective:
@@
This top-$k$ sparsity constraint has been proven to support identifying monosemantic neurons at scale [69]. Each neuron in the latent space becomes sensitive to a distinct recurring pattern across image-caption pairs ranging from object categories to physics violations, laying the foundation for feature-level analysis in subsequent stages. The activation of a neuron in the trained k-sparse autoencoder is given as:
$\mathcal{A}_{\text{neuron}}(\mathbf{e_{i}})=\mathrm{ReLU}({\mathbf{w_{\text{neuron}}}\cdot\mathbf{e_{i}}}+b_{\text{neuron}}).$ (3)
To capture a wide range of interpretable visual patterns, we train a large ensemble of sparse autoencoders (SAEs), each on a different subset of the dataset. This strategy promotes the diversity of learned visual patterns by ensuring that each SAE learns from a distinct portion of the image distribution. We train in total 200 (could be further scaled up) k-sparse autoencoders using OpenAI’s clip-vit-large-14 model as $\mathrm{E_{x}}$ and $\mathrm{E_{y}}$ [71] ($d=768$) to find a diverse set of neurons, with each SAE model having a latent space of $15,000$ neurons and $k=32$ for the top-k sparsity constraint. We have adopted hyperparameters similar to recent advances in the field of mechanistic interpretabiliey [72, 73]. In the exploration stage of this project, we have also tried using other alternative models [74] as the visual encoder but got less interpretable results. We therefore use CLIP models as their visual embeddings are aligned with semantic meaning and supports the finding of more explanable features. We subsequently combine them by collecting their weights from the linear layers, $\mathbf{w_{\text{neuron}}},b_{\text{neuron}}$.
Feature Categorization with LMMs
@@
After training the sparsity-driven interpretability modules, we obtain a collection of latent neurons, each of these neurons is activated in response to a specific feature. We employ a prompt engineering pipeline that uses large multimodal models (LMMs) [75, 76] to analyze and interpret these latent neurons.
@@
In specific, for each neuron, we collect a subpopulation [77] of image-text pairs that trigger strong activations. We then prompt Claude-3.5-Haiku[75] to summarize the common characteristics within these samples. The prompt we use to summarize the commonality of the feature is given below (more prompts in Appendix C):
@@
Feature Summarization Prompt ($\text{Prompt}_{\text{sum}}$) You are an expert in multimodal feature analysis. Given the following images and texts: {samples}, analyze the commonalities among them. Concisely summarize one common feature that is possessed by all the instances. Wrap the commonality in square brackets.
@@
Based on their returned summarization of commonalities, we categorize the features into one of the nine humanly-designed subgroups, $\mathbf{C}$. These subgroups include five categories designed for semantics: human, animal, object, activity, and environment; two categories to capture unrealistic artifacts: style and artifact; and two to capture physics violations: distortion and structure (See Appendix G). We use the following prompt to categorize the neuron based on their explanations.
@@
Feature Categorization Prompt ($\text{Prompt}_{\text{cat}}$) We have a group of datapoints possessing one common feature. Given the description of their explanation: {explanation}, which of the five categories does the feature best belong to, human, animal, object, activity, or environment? Only output the category and wrap the answer in square brackets.
Following notations in previous theoretical analysis on prompt engineering [78], the feature analysis procedures on a LanSE neuron can be formulated as below:
@@
Detailed prompts on how we use LMMs for feature analysis are listed in Appendix C. Noticing that the annotations are obtained during the construction of LanSEs, no further assistance are needed from LMMs during evaluation inferences.
@@
We denote our visual-textual LanSEs as $\mathrm{S_{\text{c}}}:(\mathcal{X},\mathcal{Y})\rightarrow\mathbb{R}^{d_{\text{c}}}$, ($\text{c}\in\mathbf{C}$). From the trained sparse auto-encoders, we get their weight matrices by concatenating neurons in corresponding neuron groups: $\mathbf{W_{\text{c}}}=[\mathbf{w_{\text{n}}}]$$\mathbf{b_{\text{c}}}=[b_{\text{n}}]$ ($n\in N_{\text{c}}$).
@@
Subsequently, for the semantic neuron groups (human, animal, object, activity and environment), we distill [79] the visual-textual LanSEs, $\mathrm{S_{\text{c}}},\text{c}\in\mathbf{C}$, into single-modality LanSEs, $\mathrm{S^{img}_{\text{c}}}$ and $\mathrm{S^{txt}_{\text{c}}}$ ($\text{c}\in\mathbf{C}$). The LanSEs are initialized with pretrained visual and textual encoders and tuned with parameter efficient fine-tuneing [80]. In specific, we tune the last layer of the CLIP model and attach a tunable head from the CLIP embedding layer to the sparse latent space. We minimize the L2-norm between the outputs of the modality-specific encoders and their corresponding activations in the joint model:
@@
This modality distillation procedure results in models that can map single-modality inputs, such as image-only or text-only, to the shared sparse latent space, enabling us to apply LanSE even when only images are available, and to analyze which features are encoded in one modality but not the other.
Conclusion
We present LanSE, a novel architecture that enables interpretable, fine-grained evaluation of generative models by indentifying interpretable visual patterns and grounding them in natural language. Built on a large pool of curated interpretable neurons, LanSE offers four actionable metrics, prompt match, visual realism, physical plausibility, and content diversity, that capture key dimensions of generation quality. Extensive human and automated evaluations validate LanSE’s reliability and alignment with human judgment. By bridging mechanistic interpretability with practical evaluation needs, LanSE contributes a robust, human-aligned framework that empowers more transparent, accountable, and controllable generative systems. Our models and evaluation suite are released to support further research and deployment.

### v2

@@ Results
Figure 9: CLIP score’s blind spots revealed through LanSE’s fine-grained analysis. Examples of images having high CLIP scores (>0.28>0.28) with obvious semantic mismatches. LanSE successfully detects mismatches that CLIP overlooks.
Figure 10: LanSE enables group-wise semantic mismatch detection beyond CLIP score. Scatter plots comparing CLIP scores with LanSE’s prompt match metric ($\mathcal{M}_{prompt}$) show agreement on severe misalignments (empty upper-right) but reveal LanSE’s unique ability to identify specific semantic failures—distinguishing whether mismatches occur in human, animal, object, activity, or environment categories.
Figure 11: LanSE reveals model differences that traditional metrics fail to capture. Evaluation results across traditional metrics (FID, CLIP, IS) show limited discriminative power and occasionally misalign with human preferences. In contrast, LanSE-derived diagnostic metrics captures these differences as shown in Figure 6.
LanSE Beats Prevalent LMMs in Fine-grained Error Detection
In this era of generative AI, many researchers consider LMMs with prompt engineering as a quick and universal solution for diverse problems [68, 69, 70]. Comparatively, LanSE can detect fine-grained visual patterns that state-of-the-art LMMs miss, while requiring significantly lower computational resources during inference. We compare LanSE’s specialized error detection capabilities against six state-of-the-art vision-language models on identifying physical plausibility errors. Using 3,410 human-annotated images (2,522 with physics violations, 888 without), we tasked each system with detecting anatomical impossibilities, distorted structures, and physics violations.
As shown in Figure 12, vision-language models showed limited performance on this task. While GPT-4o achieved 72.3% accuracy, other prominent models performed poorly: Claude-3-Haiku (34.5%), Gemini-Pro-1.5 (33.2%), Claude-3.5-Sonnet (28.7%), Qwen-2-7B (28.6%), and Claude-3-Opus (27.7%). These models consistently missed subtle anatomical errors such as extra fingers, impossible joint configurations, and localized physics violations.
LanSE achieved the best accuracy (76.6%) using 383 specialized neurons (163 for structural errors, 220 for distortions) that explicitly encode specific failure patterns. This targeted approach demonstrates that interpretable sparse representations can capture fine-grained visual anomalies that general-purpose vision-language architectures miss, highlighting the value of domain-specific pattern detection for content analysis applications.
Figure 12: LanSE outperforms LMMs in detecting physical plausibility errors. Comparison of accuracy across methods on 3,410 human-annotated images shows LanSE’s specialized neurons achieve superior performance in identifying anatomical impossibilities and structural violations that LMMs consistently underperform.
@@
We are witnessing a profound societal transformation as AI-generated content gradually permeates nearly every domain of human content creation, fundamentally challenging how we understand, verify, and trust the materials we encounter. Research across multiple domains documents serious societal risks from AI-generated content proliferation, including homogenization of human expression [20] and creative diversity [19], vulnerability to data poisoning attacks in medical systems [14], model collapse from recursive training on synthetic data [22], contamination of scientific datasets with undetected artificial material [71], and erosion of editorial standards in professional publishing [17]. These documented challenges underscore the critical need for comprehensive content analysis infrastructure that can systematically characterize AI-generated content properties, enabling informed governance and deployment of these materials across high-stakes domains. LanSE provides the systematic content analysis infrastructure this transformed landscape demands, automatically discovering thousands of interpretable visual patterns that enable comprehensive, human-understandable assessment of AI-generated content properties.
LanSE tackles these diverse challenges by combining pattern-based content analysis with distribution-level diversity assessment, enabling both granular detection of specific content characteristics and systematic measurement of variation trends. The pattern identification capability allows LanSE to detect specific synthetic signatures that indicate data contamination, identify anomalous visual patterns that suggest medical data poisoning, and recognize quality degradation markers in editorial and publishing contexts. Meanwhile, diversity assessment capabilities enable early detection of the content narrowing that precedes model collapse and systematic measurement of homogenization trends across creative outputs that threaten cultural and expressive diversity.
LanSE’s systematic pattern analysis transforms how critical institutions—from hospitals to regulatory agencies—understand and govern AI-generated content. In medical diagnostics and financial decision-making, LanSE provides clinicians and analysts with language-grounded explanations of AI predictions, delivering the transparency and accountability these domains require. Regulators can customize LanSE’s pipeline to detect domain-specific risks for educational, publishing, and safety-sensitive materials, enabling precise content filtering with transparent natural language justifications. For machine learning researchers, LanSE enables systematic dataset analysis by quantifying pattern distributions to reveal bias and guide targeted augmentation, while providing early warning signals for model collapse through diversity measurement. Beyond analysis applications, LanSE enables integration into training pipelines for interpretable model feedback and supports development of pattern-based retrieval systems where users can find content sharing specific visual characteristics. For the interpretability field, LanSE represents a paradigm shift—repositioning discovered patterns from primarily explanatory tools to functional infrastructure for diverse downstream applications.
While LanSE demonstrates versatility across multiple applications, several avenues could strengthen and extend this method. First, expanding the scale and granularity of discovered visual patterns would enhance evaluation precision, as current assessments remain bounded by our pattern vocabulary. Second, although we successfully adapted LanSE to medical imaging through CXR-LanSE, extending the framework to additional modalities, video, 3D content, or scientific imaging, presents promising opportunities. Third, while we currently demonstrate LanSE primarily for content analysis, dataset characterization, and model correlation studies, we envision integrating LanSE directly into model training pipelines, enabling generative models to receive diverse, interpretable feedback on specific visual patterns during optimization. Fourth, developing adaptive pattern vocabularies that evolve with emerging AI capabilities could maintain LanSE’s effectiveness as generative models advance.
In conclusion, LanSE establishes essential infrastructure for navigating the era of ubiquitous AI-generated content. By providing systematic, interpretable analysis of visual patterns, LanSE transforms how society understands, governs, and deploys AI-generated material across critical domains. This pattern-based content analysis approach enables diverse institutions to make informed decisions about AI integration across applications from medical diagnostics to content moderation. As AI-generated content becomes increasingly prevalent in high-stakes applications, LanSE provides the analytical foundation necessary for navigating this transformation with deliberate oversight and control.
@@
In this section, we present LanSE’s systematic content analysis pipeline (See Figure 13). Our method addresses the challenge of decomposing AI-generated content into language-grounded visual patterns through a three-stage approach: (1) discovering interpretable neurons using interpretability modules (2) categorizing these neurons automatically with LMMs, and (3) building single-modality LanSEs through a modality distillation procedure. We will release our models, scripts, and human-annotated datasets at https://github.com/YimingTangible-NUS/LanSE.
Figure 13: Pipeline for constructing Language-Grounded Sparse Encoders. The method proceeds through pattern discovery using sparse autoencoders and transcoders, automated categorization with LMMs, and modality distillation to enable single-input content analysis across diverse applications.
Dataset Construction
We use a collection of image-caption pairs from MS-COCO [72], Flickr8k and Flickr30k [73], TextCaps [74], NoCaps [75], and MIMIC-CXR [76] to build LanSE and CXR-LanSE. We augment these datasets with samples from several generative models [62, 63, 64, 65, 77, 78] by prompting them with captions from the original collections. We also employ human data annotators to identify erroneous images from the dataset for targeted discovery of physical plausibility patterns (detailed in Appendix D).
Visual Pattern Discovery with Interpretability Modules
To systematically discover interpretable visual patterns for content analysis, we leverage recent advances in mechanistic interpretability [43, 44, 45, 46, 47, 48]. Studies have shown that enforcing sparsity during neural network training produces monosemantic neurons—individual units that respond to specific, interpretable concepts [79, 80, 81, 82, 83], motivating us to adapt these techniques for content analysis.
In this work, we employ sparse autoencoders and transcoders to discover visual patterns that are interpretable in natural language, enabling systematic content analysis. We first encode image-caption pairs into multimodal embeddings by concatenating image and text representations, $\mathbf{e_{i}}=[\mathrm{E_{x}}(\mathbf{x_{i}}),\ \mathrm{E_{y}}(\mathbf{y_{i}})]$, then utilize these embeddings to train the pattern discovery modules.
Sparse autoencoders are sparsity-driven interpretability modules that can comprehensively discover a diverse set of general visual patterns through the reconstruction of one model embedding. We develop multiple sparse autoencoders, each comprising a single-layer encoder $\mathbf{z_{i}}=\mathrm{Enc}(\mathbf{e_{i}})=\mathrm{ReLU}(\mathbf{w_{enc}}\cdot\mathbf{e_{i}}+\mathbf{b_{enc}})$ and a single-layer decoder $\mathrm{Dec}(\mathbf{z_{i}})=\mathrm{ReLU}(\mathbf{w_{dec}}\cdot\mathbf{z_{i}}+\mathbf{b_{dec}})$. A top-$k$ activation scheme [45] is used to encourage sparsity in the latent space: for each $\mathbf{z_{i}}$, we retain only the $k$ highest-magnitude activations and set the rest to zero. The training objective is as follows:
@@
Transcoders are sparsity-driven interpretability modules that enable targeted discovery of specific pattern groups by learning sparse mappings from task-specific classifier representations to target outputs. We train transcoders to predict the hidden representations of lightweight classifiers designed to detect specific content properties (e.g., physics violations). This targeted approach complements sparse autoencoders by discovering rare but critical visual patterns that might be underrepresented in general reconstruction-based discovery. The training objective follows:
$\mathcal{L}_{\text{transcoder}}=\mathbb{E}_{(\mathbf{x_{i}},\mathbf{y_{i}})}\left[\left\|\mathrm{Dec}(\mathrm{TopK}(\mathrm{Enc}(\mathbf{e_{i}})))-\mathbf{t_{i}}\right\|_{2}^{2}\right].$ (2)
where $\mathbf{t_{i}}$ represents classifier hidden activations that contain task-relevant information.
The neurons discovered by these interpretability modules are initially raw and noisy, requiring systematic curation to identify reliably interpretable patterns. To maximize pattern discovery coverage while enabling quality filtering, we train multiple modules on different data subsets. In total, we train 200 sparse autoencoders and 50 transcoders for LanSE, plus 100 sparse autoencoders for CXR-LanSE, with each module containing 15,000 latent neurons and k=32 for the top-k sparsity constraint. From this large pool of candidate neurons, we apply systematic filtering criteria to identify those with consistent, interpretable activations. We adopt hyperparameters similar to recent developments in mechanistic interpretability [84, 85] and the approach can be scaled with additional compute and data.
Visual Pattern Categorization with LMMs
@@
After training the interpretability modules, we obtain a collection of latent neurons, each of them potentially interpretable in natural language. We then employ LMMs [86, 87] to analyze and interpret these latent neurons.
@@
For each neuron, we collect a set of samples that trigger strong activations for this neuron and prompt Claude-3.5-Haiku [86] (a large multimodal model) to summarize the common characteristics within these samples. The prompt we use to summarize the visual pattern in natural language is given below (more prompts in Appendix C):
@@
Pattern Summarization Prompt ($\text{Prompt}_{\text{sum}}$) You are an expert in visual pattern analysis. Given the following images and texts: {samples}, analyze the commonalities among them. Concisely summarize one common visual pattern that is possessed by all the instances. Wrap the commonality in square brackets.
@@
Based on the natural language summarizations of the visual patterns, we categorize these visual patterns into nine groups, $\mathbf{C}$, including five groups designed for semantics: human, animal, object, activity, and environment; two groups for visual realism: style and artifact; and two groups for physical plausibility: distortion and structure (See Appendix G). We use the following prompt to categorize the visual pattern based on their explanations.
@@
Pattern Categorization Prompt ($\text{Prompt}_{\text{cat}}$) We have a group of datapoints possessing one common visual pattern. Given the description of their explanation: {explanation}, which of the five categories does the visual pattern best belong to, human, animal, object, activity, or environment? Only output the category and wrap the answer in brackets.
These visual pattern analysis procedures can be formulated as below [88], noticing that the summarizations are obtained automatically during the construction of LanSEs, no assistance from LMMs are used in inference.
@@
LanSE supports two pathways for caption-free evaluation: direct training on visual-only datasets, or modality distillation from visual-textual LanSE models. For visual-textual LanSEs, we develop a modality distillation procedure as an inherent feature that enables applications to image-only datasets.
We denote one visual-textual LanSE as $\mathrm{S_{\text{c}}}:(\mathcal{X},\mathcal{Y})\rightarrow\mathbb{R}^{d_{\text{c}}}$, ($\text{c}\in\mathbf{C}$), with its language-grounded sparse layer obtained by concatenating neurons in the interpretability modules: $\mathbf{W_{\text{c}}}=[\mathbf{w_{\text{n}}}]$, $\mathbf{b_{\text{c}}}=[b_{\text{n}}]$ ($n\in N_{\text{c}}$).
@@
Modality distillation refers to training single-modality LanSEs, $\mathrm{S^{img}_{\text{c}}}$ and $\mathrm{S^{txt}_{\text{c}}}$, using outputs from $\mathrm{S_{\text{c}}}(\mathbf{x},\mathbf{y})$ [89]. The LanSEs are initialized with pretrained encoders and tuned with parameter efficient fine-tuneing [90]. The optimization target is to minimize the distances between the outputs of the single-modality LanSEs and $\mathrm{S_{\text{c}}}(\mathbf{x},\mathbf{y})$:
@@
This modality distillation procedure results in LanSEs that only require single-modality inputs, enabling us to apply LanSE when only images are available, and to analyze which patterns are encoded in one modality but not the other.